OpenStudy (anonymous):

Solve Pls Help 0=80p^2-125

OpenStudy (anonymous):

solve for p^2 first.

OpenStudy (anonymous):

ok...?

OpenStudy (anonymous):

p^2 = ? give it in fraction form.

OpenStudy (anonymous):

Idk?

OpenStudy (anonymous):

start with your equation and add 125 to both sides of the = sign

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

125 = 80p^2... now divide both sides by 80.. so you'll end up with 125/80. this is p^2. can you reduce the fraction?

OpenStudy (anonymous):

Still having trouble understanding, but Continue

OpenStudy (anonymous):

125/80 = p^2 . reducing the fraction on the left side, you'll end up with 25/16. so your equation is now 25/16 = p^2

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

wait then you square it, right?

OpenStudy (anonymous):

So it will be +-5/4 right?

OpenStudy (anonymous):

no, take the square root of 25/16

OpenStudy (anonymous):

ok, that's what you meant. be careful of the terms because squaring means something totally different from square root"ing"

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

np